OpenStudy (anonymous):

can sine or cosine ever equal 5? why or why not? I think their value can never exceed 1 but not sure why or it that is even right.

OpenStudy (amistre64):

since and cosine have a range of output values from -1 to 1

OpenStudy (amistre64):

*sine and cosine ...

OpenStudy (anonymous):

how do I support that answer

OpenStudy (amistre64):

if we alter the parent function by mutliplying it by 5, then the new function will have a range between -5 and 5, but altering a function is not quite what the question is asking tho

OpenStudy (amistre64):

depends on what your material defines sine and cosine to be

OpenStudy (amistre64):

they are by default, functions of a right triangle whose hypotenuse is 1

OpenStudy (anonymous):

what does that mean with the right triangle

OpenStudy (amistre64):

1^2 = x^2 + y^2 as such, y = sqrt(1-x^2) since a sqrt has no real value for negative numbers: 1-x^2 has to be 0 or positive 1 - x^2 >= 0 1 >= x^2 1 >= x

OpenStudy (amistre64):

sine and cosine are functions that relate the ratio of the leg of a right triangle to the hypotenuse.

OpenStudy (anonymous):

so, I don't quite undertstand the reas why they can never =5

OpenStudy (amistre64):

if something is either red or blue, can it ever be green?

OpenStudy (anonymous):

oh, because of sine =opp/hyp and cos=adj/hyp?

OpenStudy (amistre64):

sine and cosine only have outputs from -1 and 1, as such they can never be 5 since 5 is not between -1 and 1

OpenStudy (amistre64):

yes

OpenStudy (amistre64):